Find the Cartesian equation of the locus of the point P representing the complex
number z. Sketch the locus of P in each case.
(a) 2|z + 1| = |z − 2|
1
Expert's answer
2020-05-26T18:29:11-0400

"2|z+1|=|z-2|"

"2|x+iy+1|=|x+iy-2|"

"2^2|(x+1)+iy|^2=|(x-2)+iy|^2"

"4((x+1)^2+y^2)=(x-2)^2+y^2"

"4x^2+8x+4+4y^2=x^2-4x+4+y^2"

"3x^2+12x+3y^2=0"

"x^2+4x+y^2=0"

"(x^2+4x+4)+y^2=4"

"(x+2)^2+y^2=4"


Answer: the locus is a circle with center "(-2,0)" and radius "2" .